# Distribution-aware Knowledge Unification and Association for Non-exemplar Lifelong Person Re-identification

### v1

@@ IV-E Ablation Studies
To assess the contributions of individual components in our DKUA, we conduct ablation studies to determine the impact of Baseline, knowledge alignment (KA), unified knowledge association (UKA), distribution-based knowledge transfer (DKT) on model performance. Baseline consists of distribution-aware model and knowledge unification. These experiments are performed on unseen and seen domains under training order-1, as illustrated in Table III.
Performance of Different Components: In Table III, the Baseline demonstrates that our distribution-aware model and knowledge unification strategy effectively capture domain-specific distribution and consolidate unified representation across all domains for alleviating catastrophic forgetting. Compared to the second row, our KA enforces narrowing the distance between domains, ensuring that unified representation learning remains accurate and effective. Compared to the second and third row, our unified knowledge association mechanism achieves a significant improvement of 3.1% mAP/4.8% R@1, and 3.5% mAP/4.1% R@1. This demonstrates that our UKA effectively associates knowledge between domain-specific and unified representations based on each instance to reduce inter-domain gaps for enhancing the model’s anti-forgetting and generalization. Compared to the third and fourth rows, our DKT design enables the model to transfer domain-specific distributions into unified knowledge based on all exemplars of each domain. Each independent component in our DKUA plays a crucial role. When the four components are merged, our DKUA significantly improves both the anti-forgetting and generalization capacity of the model.
@@
V Visualization
To further validate the effectiveness of our DKUA, we conduct a qualitative comparison with the DASK method, as shown in Figure. 6. The left column displays challenging query images, including cases with low-light, occlusion, back views, and side views. In Figure. 6, correctly matched pedestrians are marked with green boxes, while incorrect matches are highlighted in red. Under low-light conditions (first row), our DKUA accurately retrieves eight out of ten individuals with the same identity, despite the limited visual information available. Under occlusion conditions (second and third rows), the DASK method matches at least four people incorrectly among ten individuals. But our DKUA accurately retrieves nine out of ten individuals with the same identity. In back-view scenarios (fourth and fifth rows), our DKUA successfully handles angle change issue, outperforming DASK method in matching the same identity. For side-view (sixth row), the DASK method only correctly matches four pedestrians, whereas our DKUA achieves perfect retrieval (10/10) pedestrians with the occlusion and side view. These results demonstrate that DKUA achieves superior retrieval performance, consistently identifying more pedestrians with the same identity across diverse challenging conditions. This advantage stems from our DKUA, which effectively consolidates the unified representation with the characteristics of cross-domain sharing and identity discrimination to enhance the model’s anti-forgetting and matching accuracy performance.
VI Conclusions

### v2

@@ IV-E Ablation Studies
To assess the contributions of individual components in our DKUA, we conduct ablation studies to determine the impact of Baseline, adaptive knowledge consolidation (AKC), unified knowledge association (UKA), distribution-based knowledge transfer (DKT) on model performance. Baseline consists of distribution-aware model and knowledge unification. These experiments are performed on unseen and seen domains under training order-1, as illustrated in Table IV.
Performance of Different Components: In Table IV, the Baseline demonstrates that our distribution-aware model and knowledge unification strategy effectively capture domain-specific distribution and consolidate unified representation across all domains for alleviating catastrophic forgetting. Compared to the second row, our AKC enforces narrowing the distance between domains, ensuring that unified representation learning remains accurate and effective. Compared to the second and third row, our unified knowledge association mechanism achieves a significant improvement of 3.1% mAP/4.8% R@1, and 3.5% mAP/4.1% R@1. This demonstrates that our UKA effectively associates knowledge between domain-specific and unified representations based on each instance to reduce inter-domain gaps for enhancing the model’s anti-forgetting and generalization. Compared to the third and fourth rows, our DKT design enables the model to transfer domain-specific distributions into unified knowledge based on all exemplars of each domain. Each independent component in our DKUA plays a crucial role. When the four components are merged, our DKUA significantly improves both the anti-forgetting and generalization capacity of the model.
@@
TABLE V: Experimental results regarding training time, Parameters, and FLOPs.
Method Training time (h-min) Parameters(M) FLOPs(G)
DRE[40] 20h-18min 99.83 21.2
PAEMA[7] 6h-20min 86.03 11.3
DCR [24] 19h-10min 129.5 12.6
DKUA-ImageNet 6h-35min 102.7 18.3
DKUA 3h-22min 39.67 8.1
IV-F Analysis of Computational Complexity
Complexity Analysis of transfer module (TM): Each domain is equipped with a dedicated Transfer Module (TM) to capture its domain-specific distribution, thereby mitigating catastrophic forgetting. To quantify the increase in model complexity, we compute the FLOPs and parameters of each TM. The TM adopts the structure of a final transformer layer, which includes multi-head self-attention (MHSA), a feed-forward network (FFN), residual connections, and layer normalization. A single TM introduces 228.7 M FLOPs and 1.774 M parameters, which uses a small amount of memory compared to storing old exemplars [40]. This design allows our DKUA model to effectively learn and retain domain-specific knowledge, successfully alleviating catastrophic forgetting.
Efficiency of Computational Complexity: We conduct experiments to evaluate computational complexity in comparison with ViT-based methods, including DRE [40], PAEMA [7], DCR [24], and DKUA-ImageNet, as summarized in Table V. Training time is reported in hours (h) and minutes (min). Although our DKUA-ImageNet produces higher FLOPs and parameter counts than DRE, DCR, and PAEMA, it achieves a shorter training time than DRE and DCR. More importantly, DKUA-ImageNet delivers significantly better performance, with improvements of at least 1.5% mAP/0.7% R-1 and 2.1% mAP/2.6% R-1 on Seen-Avg, and at least 1.8% mAP/2.2% R-1 and 2.6% mAP/2.2% R-1 on Unseen-Avg for training order-1 and order-2, as shown in Tables II and III. To reduce the model’s parameters and FLOPs, our DKUA introduces a ViT-Small/16 [41] architecture as the backbone. Our strategies (such as AKC, UKA, and DKT) with a rehearsal-free and distillation-free scheme are well-suited to the backbone, significantly enhancing the model’s anti-forgetting and generalization abilities in Tables II and III.
IV-G Visualization
To further validate the effectiveness of our DKUA, we conduct a qualitative comparison with the DASK method, as shown in Fig 6. The left column displays challenging query images, including cases with low-light conditions, occlusion, back and side views. In Fig. 6, correctly matched pedestrians are marked with green boxes, while incorrect matches are highlighted in red. Under low-light conditions (first row), our DKUA accurately retrieves eight out of ten individuals with the same identity, despite the limited visual information available. In the case of occlusion (second and third rows), the DASK method produces at least four incorrect matches out of ten individuals, whereas DKUA correctly retrieves nine. For back and side views (fourth and fifth rows), our method successfully handles the perspective change issue and outperforms DASK in identifying the same identity. In the side-view scenario with occlusion (sixth row), DASK correctly matches only four pedestrians, while DKUA achieves perfect retrieval (10/10). These results demonstrate that DKUA achieves superior retrieval performance by consistently identifying more pedestrians of the same identity across diverse challenging conditions. This advantage stems from its ability to effectively learn a unified representation that captures cross-domain shared features while maintaining identity discriminability, thereby enhancing the model’s anti-forgetting capability and matching accuracy.
V Conclusions